OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7)? A. 2 B. 3/4 C. -4/3 D. -2

OpenStudy (black_ninja123):

Slope formula \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (black_ninja123):

\[\frac{ (-7)-3 }{ 4-(-1) }\]

OpenStudy (black_ninja123):

\[\frac{ -10 }{ 5 }\]

OpenStudy (black_ninja123):

D is your answer -2